A collection of quarters and nickels is worth $1.25. There are 13 coins in all. How many of each are there?
Answer by jorel1380(3719)

n+q=13
5n+25q=125
5n+5q=65
20q=60
q=3
n=10
The collection has 10 nickels and 3 quarters..